OpenStudy (anonymous):

A sector of a circle has perimeter 7cm and area 3cm^2. Find all possible radii.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

best answer

OpenStudy (superdavesuper):

lets assume the radius is r and the enclosed angle of the sector is theta, we can set up the eqns for perimeter n area....

OpenStudy (superdavesuper):



OpenStudy (ranga):

Length of the arc = r * theta (theta is in radians) Perimeter of the sector = 2r + r * theta = 7 ---- (1) Area of the sector = 1/2 * r^2 * theta = 3 ----- (2) Eliminate theta from (1) and (2) and solve for r.

OpenStudy (anonymous):

I attempted to do it but my answers are still incorrect. I keep ending up with 1.11 instead of 1.5.

OpenStudy (superdavesuper):

Nope 1.5 is one of the right answer; plz try to solve the two equations again as ranga has set up above.

OpenStudy (anonymous):

The answer is in the attachment.

OpenStudy (anonymous):

Thank you for including how to do it step by step! I was lost half way through the equation and did not know where my mistake was.